OpenStudy (anonymous):

Solve the triangle. Round lengths to the nearest tenth and angle measures to the nearest degree. a = 6, c = 11, B = 109°

OpenStudy (akashdeepdeb):

Do you know the law of sines and cosines?

OpenStudy (noorik.):

maybe you should draw the triangle

OpenStudy (akashdeepdeb):

@kittiesnkisses ?

OpenStudy (anonymous):

I don't know the law

OpenStudy (anonymous):

@AkashdeepDeb

OpenStudy (noorik.):

the law is \[\frac{ A }{ sinA} = \frac{ B }{ sinB}\]

OpenStudy (anonymous):

Where does C come in?

OpenStudy (noorik.):

also \[\frac{ B }{ sinB} = \frac{ C }{ sinC }\]

OpenStudy (anonymous):

Ohh gotcha

OpenStudy (noorik.):

hope I helped

OpenStudy (anonymous):

You did. Thank you!